OpenStudy (rfmn-dfmn):

If f(x)=2x2(squared) -10 ,find f(5)

OpenStudy (rfmn-dfmn):

A:90 B:10 C:7.5 D:40

OpenStudy (peachpi):

plug in 5 for x f(5) = 2(5)² =

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (rfmn-dfmn):

So the answer would be 10? @Peachpi

OpenStudy (peachpi):

no. use PEMDAS. square the 5 first, then multiply by 2, then subtract 10 f(5) = 2(5)² - 10 = 2(25) - 10 = 50 - 10 = 40

OpenStudy (rfmn-dfmn):

I'm sorry about all the questions but how did you get 25 @peachpi

OpenStudy (peachpi):

because 5² = 5 * 5 = 25

OpenStudy (rfmn-dfmn):

Thank you so much

OpenStudy (rfmn-dfmn):

@peachpi